### v1

# Josephson effect with periodic order parameter

Abstract:
We investigate the Josephson effect in a two-dimensional superconducting system with a smoothly and periodically varying order parameter. The order parameter is modulated along one direction while remaining uniform in the perpendicular direction, leading to a spatially periodic superconducting phase. We show that the periodicity of the order parameter determines the winding number of the eigenfunctions, which serves as a topological characterization of the system. The winding number is calculated analytically and visualized through the trajectory of the corresponding three-dimensional Bloch vector. By solving the Bogoliubov–de Gennes equation, we obtain both plane-wave solutions describing bulk states and exponentially localized solutions that correspond to edge modes. The analytic bulk–edge connection is employed to identify the conditions under which the edge states emerge from the bulk spectrum. We find that the winding numbers depend on the boundary conditions, which differ between the plane-wave and exponential solutions. These results establish a direct connection between the spatial modulation of the order parameter, the topological structure of the eigenstates, and the emergence of edge modes in periodically modulated Josephson systems.
1 Introduction
The $SU(2)$ Hamiltonian $\vec{h}\cdot\vec{\sigma}$ has been used as a protype for topological states in condensed matter to describe a number of fascinating properties in special materials, ranging from 2D graphene-like materials to 3D Weyl semimetals. Many of the specific properties originate in the robust winding number of the spinor states. A deeper understanding of the connection between the $SU(2)$ Hamiltonian and the winding number of the spinor will reveal robust physical properties of systems that are governed by such a Hamiltonian. In the following we will study conditions that can be understood as the Josephson effect in the presence of a periodically changing order parameter. The Josephson effect is typically described by two superconductors, which are separated by a small non-superconducting barrier, the Josephson junction [1, 2, 3]. The corresponding order parameters have distinct phase factors that leads to the Josephson effect that is characterized by the Josephson current, which flows between the two superconductors. Such a discontinuous behavior of the order parameter is replaced subsequently by a smoothly varying phase of the order parameter. Although this agrees only vaguely with the original concept of a Josephson junction, we find nevertheless some interesting effect in terms of the wave function properties. In particular, We will discuss that the winding number is directly determined by the order parameter phase.
The winding number is often associated with a robust behavior due to its connection with topological invariants, for instance, in chiral quantum systems. Therefore, a weak perturbation might be insufficient for a change but we must rely on a strong and macroscopic intervention [4]. What determines the winding number in a given system and how can we change it? And is it possible to control it by an external and macroscopic method? To answer these questions, we consider a torus or an open cylinder made of a superconducting material and apply a phase modulated order parameter. We will discuss that the latter is connected with the boundary conditions of the system. Alternatively, we will study a tight-binding model whose Brillouin zone is compact, for instance in the form of a torus.
2 Model
There is an intimate connection between spinors and winding numbers, which will be central for the following discussion. As an instructive demonstration of this connection, we consider the $2\times 2$ Hamiltonian
[…]
with $m$ real , $z=z^{\prime}-iz^{\prime\prime}$ and real $^{\prime}, z′′z^{\prime\prime}$. The vector $\vec{\sigma}=(\sigma_{1},\sigma_{2},\sigma_{3})$ comprises the Pauli matrices and $\vec{h}_{0}=(z^{\prime},z^{\prime\prime},m)$. The corresponding eigenvalue problem with eigenvalue $E=\pm\sqrt{m^{2}+|z|^{2}}$ reads
[…]
which separates into two equations
$\cases{a_{1}=\frac{z}{E-m}a_{2}\cr a_{2}=\frac{z^{*}}{E+m}a_{1}\cr}.$
A solution of these equations is the spinor
[…]
The winding number is observable and usually related to the Berry phase, where the latter can be calculated from the Berry connection [5, 6]. Alternatively, the spinor can also be associated with the three-dimensional Bloch vector
$\vec{s}=\frac{\Psi\cdot\vec{\sigma}\Psi}{\Psi\cdot\Psi},$ (3)
which is the expectation value of the vector $\vec{\sigma}$. The Bloch vector is of unit length and its components read
$s_{1}=\frac{2z^{\prime}}{{\cal N}_{0}(E+m)},\ \ s_{2}=\frac{2z^{\prime\prime}}{{\cal N}_{0}(E+m)},\ \ s_{3}=\frac{1}{{\cal N}_{0}}\left[1-\frac{|z|^{2}}{(E+m)^{2}}\right]$ (4)
with the normalization ${\cal N}_{0}=1+|z|^{2}/(E+m)^{2}$. This means that a closed trajectory of $z$ on the complex plane or on some Riemann sheets maps onto a closed trajectory on the Bloch sphere.
After this brief preliminary discussion we turn to the $SU(2)$ Hamiltonian $\vec{h}\cdot\vec{\sigma}$ with the three-dimensional vector
$\vec{h}=(\Delta^{\prime},\Delta^{\prime\prime},D),$ (5)
where each vector component could be a self-adjoint differential operator. For the following study, however, we focus on the case where only $D$ is a translation-invariant differential or difference operator and $\Delta^{\prime}=|\Delta|\cos(x/L)$ and $\Delta^{\prime\prime}=|\Delta|\sin(x/L)$ are the real and imaginary parts of the order parameter $\Delta$, respectively, where $|\Delta|$ is spatially uniform. To create a matrix of the form of $H_{0}$ in Eq. (1) we assume that there is a pair of complex numbers $\gamma_{j}$ ($j=1,2$) with $De^{\gamma_{j}x}=d(\gamma_{j})e^{\gamma_{j}x}$ (i.e., $e^{\gamma_{j}x}$ is eigenvector of $D$ with an eigenvalue $d(\gamma_{j})$) we introduce the spinor
[…]
to formulate the energy eigenvalue problem $\vec{h}\cdot\vec{\sigma}\Psi_{x}=E\Psi_{x}$ in matrix notation as
[…]
for real $E$. Since the order parameter $\Delta$ is periodic in $x$ with the period $2\pi L$, it reduces the translation-invariance but still enables periodic eigenfunctions. The equation is satisfied for all $x$ when $\gamma_{2}=\gamma_{1}+i/L$ and
$\frac{a_{2}}{a_{1}}=\frac{E-d(\gamma_{1})}{|\Delta|},\ \ \ \frac{a_{1}}{a_{2}}=\frac{E+d(\gamma_{2})}{|\Delta|}.$ (8)
After dropping the index of $\gamma_{1}$ (i.e., $\gamma\equiv\gamma_{1}$) this implies the quadratic equation in $E$
$E-d(\gamma)][E+d(\gamma+i/L)]=|\Delta|^{2},$ (9)
which gives the dispersions
$E_{\pm}(\gamma)=\frac{d(\gamma)-d(\gamma+i/L)}{2}\pm\frac{1}{2}\sqrt{[d(\gamma)+d(\gamma+i/L)]^{2}+4|\Delta|^{2}}$ (10)
as the solutions. The values of $\gamma$ are restricted such that $E_{\pm}(\gamma)$ is real. Then the spinor in Eq. (6) becomes
[…]
where $\alpha$ is the phase of $a_{2}/a_{1}$. Next we analyze the winding number of the spinor. First, we note that the spinor without the exponential factor is parametrized by the coordinate $x$ and it is invariant under $x\to x+4n\pi L$ for an integer $n$. Then the winding of the spinor can be associated with the three-dimensional Bloch vector of Eq. (3). Its components characterize the algebraic relations between the two spinor components of the solution $\Psi_{x}$ of Eq. (11):
$s_{1}=\frac{2b}{1+b^{2}}\cos(x/L+\alpha),\ \ s_{2}=\frac{2b}{1+b^{2}}\sin(x/L+\alpha),\ \ s_{3}=\frac{1-b^{2}}{1+b^{2}},$ (12)
where $b=|a_{2}/a_{1}|=|E-d(\gamma)|/|\Delta|$. The Bloch vector is invariant under $x\to x+2n\pi L$ for an integer $n$, which is half of the periodicity of the spinor. A closed trajectory along the $x$-direction of a torus with length $l$ maps onto the Bloch sphere as a trajectory with constant latitude that is determined by $s_{3}$ and with the winding number $w=l/2\pi L$. Thus, the winding number depends only on the order parameter phase, while the radius of the trajectory $2b/(1+b^{2})$ depends on $\gamma$ and $E$. Boundary conditions fix the parameter $L$. Here we assume periodic boundary conditions, such that we get $L=l/2\pi n$ with integer $n$, which implies $w=n$. This means that we can determine the integer winding number by choosing the phase of the order parameter $x/L$. For $E=d(\gamma)$, where $b=0$, the Bloch vector is on the north pole of the Bloch sphere, while for $b=1$ it is on the equator, and it moves to the southern hemisphere for $b>1$.
Finally, we note that the unitary transformation
[…]
creates the phase factors for the order parameter by a simultaneous transformation of the operator $D$. In other words, this transforms the real order parameter $|\Delta|$ of the Hamiltonian to a complex order parameter $\Delta$, which is accompanied by the creation of the self-adjoint operator $\tilde{D}$, whose eigenvalues are the same as those of $D$. Thus, the eigenfunctions and the structure of the spinor as well as the Bloch vector and its winding number are transformed. In Sect. 3.1 we will discuss how such a unitary transformation can be realized in a physical system.
Many physical quantities depend only on the spectral properties of the Hamiltonian. For instance, trace expressions such as the thermodynamic quantities or the density of states. This is also the case for expressions represented by an inner product. These quantities are invariant under unitary transformations. Correlation functions, on the other hand, may not be invariant under general unitary transformations, such as the relation of the spinor components in Eqs. (6) or (11), whose real and imaginary parts are observable.
3 Examples
In the following we will consider two examples for the operator $D$, namely the Bogoliubov de Gennes (BdG) Hamiltonian in a continuous space and its tight-binding version on a lattice. These examples are complemented with the 2D Dirac Hamiltonian in the continuum and on the lattice.
The BdG Hamiltonians with different symmetries have been intensively studied for translation-invariant systems  [7, 8, 9]. We will rely here on its simplest form but break the translational invariance by a periodic order parameter $\Delta$, as introduced in the previous section.
3.1 Continuous BdG Hamiltonian
We choose the $SU(2)$ Hamiltonian $\vec{h}_{BdG}\cdot\vec{\sigma}$ with $D=-\partial_{x}^{2}-\partial_{y}^{2}$ acting on a finite continuous space:
$\vec{h}_{BdG}=(\Delta^{\prime},\Delta^{\prime\prime},-\partial_{x}^{2}-\partial_{y}^{2}).$ (14)
When we assume that $\Delta$ is uniform in the $y$ direction, we can apply the Fourier transformation $\partial_{y}^{2}\to k_{y}^{2}$. The corresponding eigenvalue problem of the quasiparticles with energy $E$ reads as the BdG equation
[…]
whose solutions describe the wavefunctions along a torus or a cylinder in the $x$ direction, respectively. For the BdG Hamiltonian a spatial variation of the phase induces a supercurrent ${\bf j}_{s}$ and vice versa, based on the Ginzburg-Landau supercurrent-phase relation
${\bf j}_{\rm s}=\frac{\hbar e^{*}}{m^{*}}|\Delta|^{2}(\nabla\varphi-\frac{e^{*}}{\hbar c}{\bf A})$ (16)
in the presence of a vector potential $\bf A$}. $e^{*}$ and $m^{*}$ are the charge and the mass of the Cooper pairs. This relation provides the unitary transformation in Eq. (13) by an appropriate choice of the supercurrent and the vector potential. In particular, the special choice $\partial_{x}\varphi=1/L$ creates the phase of Eq. (15).
Figure 1: The Riemann surface of the energy in Eq. (18) with complex $\gamma$ is symmetric with respect to ${\rm Re}\gamma$.
From Sect. 2 we get for this special case the $\gamma\to-\gamma$ symmetric eigenvalues $d(\gamma)=-\gamma^{2}+k_{y}^{2}$. Then Eq. (9) becomes
$E+\gamma^{2}-k_{y}^{2}][E-(\gamma+i/L)^{2}+k_{y}^{2}]=|\Delta|^{2}.$ (17)
Moreover, considering a constant solution in $y$ direction (i.e., $k_{y}=0$) we obtain as the solution of the quadratic equation $(E+\gamma^{2})[E-(\gamma+i/L)^{2}]=|\Delta|^{2}$ for $E$ the dispersion
$E_{\pm}(\gamma)=\frac{i\gamma}{L}-\frac{1}{2L^{2}}\pm\frac{1}{2}\sqrt{[\gamma^{2}+(\gamma+i/L)^{2}]^{2}+4|\Delta|^{2}}$ (18)
that is not symmetric under $\gamma\to-\gamma$ but its real part is symmetric with respect to ${\rm Re}\gamma\to-{\rm Re}\gamma$ (cf. Fig. 1). With $\gamma=\kappa+ik$, where $k$ and $\kappa$ are real, we can distinguish plane-wave solutions for $\kappa=0$ and exponential solutions for $\kappa\neq 0$. Exponential means here that the absolute value of the solution either decreases or increases exponentially with $x$. Employing the plane-wave solution $\gamma=ik$ we obtain $(E-k^{2})[E+(k+1/L)^{2}]=|\Delta|^{2}$. This gives for the dispersion in Eq. (10)
$E_{\pm}(k)=-\frac{1}{2L^{2}}-\frac{k}{L}\pm\frac{1}{2}\sqrt{[k^{2}+(k+1/L)^{2}]^{2}+4|\Delta|^{2}},$ (19)
which is real for any $k$ with $\infty<k<\infty$. This dispersion is plotted in Fig. 2a) for $|\Delta|=L=1$. Expansion for small $k$ yields a gap and a linear $k$ term:
$E_{\pm}(k)=\pm\frac{\sqrt{4L^{4}|\Delta|^{2}+1}-1}{2L^{2}}+\frac{4L^{4}|\Delta|^{2}\mp\sqrt{4L^{4}|\Delta|^{2}+1}+1}{4L^{5}|\Delta|^{2}+L}k+O(k^{2}),$ (20)
while the asymptotic behavior for large $k$ is parabolic:
$E_{\pm}\sim\pm k^{2}.$ (21)
For the zero mode $E=0$ and $k_{y}=0$ we get directly from Eq. (17) the quartic equation $\gamma^{2}(\gamma+i/L)^{2}=-|\Delta|^{2}$ that reduces to the two quadratic equations $\gamma(\gamma+i/L)=\pm i|\Delta|$ with the four solutions
$\gamma=-\frac{i}{2L}\pm i\sqrt{i|\Delta|+1/4L^{2}}\ \ {\rm and}\ \ \gamma^{\prime}=-\frac{i}{2L}\pm i\sqrt{-i|\Delta|+1/4L^{2}}.$ (22)
These solutions are exponential for $|\Delta|>0$. There are other exponential solutions for $E\neq 0$, as indicated in Figs. 3 a), b).
The unitary transformation of Eq. (13) reads in this case
$\tilde{D}=(\partial_{x}-i/2L)^{2},$ (23)
where $i/2L$ represents a gauge field.
Finally, for the plan-wave solutions we have $b=|E-k^{2}|/|\Delta|$, such that the Bloch vector of Eq. (12) reads
$s_{1}=\frac{2|E-k^{2}||\Delta|}{|\Delta|^{2}+(E-k^{2})^{2}}\cos(x/L),\ \ s_{2}=\frac{2|E-k^{2}||\Delta|}{|\Delta|^{2}+(E-k^{2})^{2}}\sin(x/L),\ \ s_{3}=\frac{|\Delta|^{2}-(E-k^{2})^{2}}{|\Delta|^{2}+(E-k^{2})^{2}}.$ (24)
It hits the north pole of the Bloch sphere for $k^{2}=E$.
a) Refer to caption b) Refer to caption
Figure 2: Energy dispersion $E_{\pm}(k)$ of the plane-wave solution for a) the continuous BdG Hamiltonian and b) for the tight-binding BdG Hamiltonian with $|\Delta|=L=1$. These dispersions correspond to the vertical lines in Fig. 2.
Figure 3: Real-E spectrum: Curves with real energy after analytic continuation of $\gamma$ with $|\Delta|=L=1$. Continuum model: a) $E_{+}(\gamma)$ and b) $E_{-}(\gamma)$. Tight-binding model: c) $E_{+}(\gamma)$ and d) $E_{-}(\gamma)$. The vertical lines with ${\rm Re}\gamma=0$ represent the plane-wave solutions, while the other curves represent exponential solutions. The corresponding dispersions for the plane waves are plotted in Fig. 2.
Figure 4: $(s_{1},s_{2})$ trajectories of the Bloch vector for the $\pi$-flux Hamiltonian with $m=1$ and different values of the radius $K=\sqrt{k_{x}^{2}+k_{y}^{2}}$ in the Fourier space.
3.1.1 2D Dirac Hamiltonian
By comparing the BdG Hamiltonian in Eq. (14) with the massive 2D Dirac Hamiltonian $H_{Dirac}=\vec{h}_{Dirac}\cdot\vec{\sigma}$ with $\vec{h}_{Dirac}=(k_{x},k_{y},m)$ we get with the 2D Dirac spinor
[…]
with the eigenvalues $E=\pm\sqrt{m^{2}+k^{2}}$ and with the Bloch vector components
$s_{1}=-\frac{k_{x}}{E},\ \ s_{2}=-\frac{k_{y}}{E},\ \ s_{3}=-\frac{m}{E}.$ (26)
This can also be written as the relation $\vec{s}_{Dirac}=-\vec{h}_{Dirac}/E$. Thus, the winding number for a closed trajectory in $k_{x}-k_{y}$ space around $k_{x}=k_{y}=0$ is $w_{Dirac}=1$. This comparison reflects an advantage of the BdG Hamiltonian with complex order parameter in place of a real mass of the 2D Dirac particle, since it enables us to change and control the winding number. However, a special form of the unitary transformation in Eq. (13) can be applied to the 2D Dirac Hamiltonian with $\vec{h}_{D}=(i\partial_{x},i\partial_{y},m)$ as $H_{D}\to H_{D}^{\prime}=U_{D}H_{D}U_{D}^{\dagger}$ with
[…]
With $\partial_{y}\varphi=0$ and after a Fourier transformation $i\partial_{y}\to k_{y}$ we can set $k_{y}=0$ and obtain $H_{D}^{\prime}=-(\partial_{x}\varphi/2)\sigma_{0}+\vec{h}_{D}^{\prime}\cdot\vec{\sigma}$ with $\vec{h}_{D}^{\prime}=(m\cos\varphi,m\sin\varphi,i\partial_{x})$, where $\sigma_{0}$ is the $2\times 2$ unit matrix. Thus, the unitary transformation creates a phase factor to the Dirac mass $m\to e^{i\varphi}m$ and an energy shift $(\partial_{x}\varphi/2)$. This yields for the Bloch vector the same form as in Eq. (12). A similar unitary mapping was studied for a double-layered chiral superconductor with circular symmetry onto two Dirac models with opposite mass signs [10]. Thus, a unitary transformation creates a phase factor for the the Dirac mass, such that this transformation implies a change of the winding number. In more practical terms, a physical system is modeled by a specific Hamiltonian, where a unitary transformation yields a different physical system with the same spectral properties.
For a lattice version of the 2D Dirac Hamiltonian ($\pi$-flux model [11]) we return to $H_{0}$ in Eq. (1) with $z=\sin k_{x}-i\sin k_{y}$ and for the Bloch vector in Eq. (4)
$s_{1}=\frac{2(E+m)}{\sin^{2}k_{x}+\sin^{2}k_{y}+(E+m)^{2}}\sin k_{x},\ \ s_{2}=\frac{2(E+m)}{\sin^{2}k_{x}+\sin^{2}k_{y}+(E+m)^{2}}\sin k_{y},$
$s_{3}=\frac{\sin^{2}k_{x}+\sin^{2}k_{y}-(E+m)^{2}}{\sin^{2}k_{x}+\sin^{2}k_{y}+(E+m)^{2}}$ (28)
with $E=\pm\sqrt{m^{2}+\sin^{2}k_{x}+\sin^{2}k_{y}}$. There are four spectral nodes at $k_{j}=0,\pi$, where each of these nodes contributes either with positive or with negative chirality. This results in more complex trajectories on the Bloch sphere for larger values of $K=\sqrt{k_{x}^{2}+k_{y}^{2}}$, as illustrated in Fig. 4, while smaller values (e.g., $K=1$) yield trajectories similar to the circles of the continuous Dirac Hamiltonian. This reflects the well-known effect of the lattice structure on topological properties [11].
3.2 Tight-binding BdG Hamiltonian
After we have seen that the winding trajectory of the Bloch vector is disturbed by the discrete lattice structure in the case of the 2D Dirac Hamiltonian, we will study next whether this is also the case for the lattice BdG Hamiltonian. To this end we consider the case in which $D$ is a difference operator with $D\psi_{x}=\psi_{x+1}+\psi_{x-1}-2\psi_{x}$ that acts on a discrete lattice with unit lattice spacing. Its eigenvalue condition reads in this case
$De^{\gamma_{j}x}=e^{\gamma_{j}(x+1)}+e^{\gamma_{j}(x-1)}-2e^{\gamma_{j}x}=2(\cosh\gamma_{j}-1)e^{\gamma_{j}x},$ (29)
such that $d(\gamma)=2(\cosh\gamma-1)$ is symmetric with respect to $\gamma\to-\gamma$ and gives with Eq. (10) the dispersion
$E_{\pm}(\gamma)=\cosh(\gamma)-\cosh(\gamma+i/L)\pm\sqrt{[\cosh(\gamma)+\cosh(\gamma+i/L)-2]^{2}+|\Delta|^{2}},$ (30)
which is complex for general $\gamma$ and not symmetric. For $\gamma=ik$ the plane-wave dispersion follows as
$E_{\pm}(k)=\cos k-\cos(k+1/L)\pm\sqrt{[\cos(k)+\cos(k+1/L)-2]^{2}+|\Delta|^{2}},$ (31)
which is real again for any $k$ with $\infty<k<\infty$. This dispersion is visualized in Fig. 1b). There are also exponential solutions with a real dispersion. They have nonzero ${\rm Re}\gamma$ values and are indicated as horizontal curves in Figs. 2 c) and d).
Inserting the eigenvalues $d(\gamma)$ into the spinor of Eq. (11) and into the Bloch vector of Eq. (12) enables us to calculate the winding number for the tight-binding BdG Hamiltonian. In contrast to the lattice effect in the Dirac Hamiltonian, the trajectories of the Bloch vector and the winding number agree with those of the continuous BdG Hamiltonian. Finally, the unitary transformation of Eq. (13) reads in this case
$\tilde{D}\psi_{x}=e^{-i/2L}\psi_{x+1}+e^{i/2L}\psi_{x-1}-2\psi_{x},$ (32)
where $i/2L$ represents a Peierls phase on the hopping elements in $x$ direction, in analogy with the gauge transformation of the continuous BdG Hamiltonian in Eq. (23).
4 Exponential solutions
To find the exponential solutions, we can either solve equation (9) directly or employ the analytic bulk-edge connection [10, 12]. The eigenvalues $d(\gamma)=-\gamma^{2}+k_{y}^{2}$ of the continuous BdG Hamiltonian and $d(\gamma)=2(\cosh\gamma-1)$ of the tight-binding BdG Hamiltonian are both real for real as well as purely imaginary $\gamma$, while the dispersions in Eq. (18) and in Eq. (30) are real only for purely imaginary $\gamma$. This suggests that the analytic continuation starts from a plane-wave (bulk) solution. Then the analytic continuation $ik\to\gamma$ of the real wave number $k$ into the complex plane is applied. In general, this leads to a complex energy, which violates the eigenvalue equation for real energies. Therefore, the additional condition of a real $E_{\pm}(\gamma)$ must be enforced. The results of the analytic bulk-edge connection are visualized in Figs. 2 a) – d): For the continuous BdG Hamiltonian we obtain the real-E curves in Figs. 2 a), b), while the tight-binding BdG Hamiltonian gives Figs. 2 c),d). In both cases the vertical lines at ${\rm Re}\gamma=0$ represent the plane-wave solutions, while the other curves represent exponential solutions. The periodic behavior of the tight-binding spectrum in Eq. (31) for imaginary $\gamma$ is reflected by the repeated real-E curves in Figs. 2c), d). Apart from this feature, the real-E spectrum is similar for both Hamiltonians and their real-E curves are symmetric with respect to ${\rm Re}\gamma\to-{\rm Re}\gamma$.
The Bloch vector in Eq. (12) is valid for plane-wave as well as for exponential solutions of Eq. (15), an exponential solution might have a phase shift $\alpha\neq 0$ and a different $b$ though. The winding number is directly linked to the phase of the order parameter $x/L$. In general, we might consider non-periodic boundary conditions, independent of the value of $L$. This indicates that the winding number of the bulk modes differ from those of the edge modes simply due to different boundary conditions. The latter depend on the specific set-up of the physical system. It is beyond the scope of the present work to elaborate on this experiment-specific issue. In any case, though, the boundary conditions must also lead to a self-adjoint Hamiltonian, which requires that the energy eigenstates for different eigenvalues are orthogonal.
5 Discussion and summary
The mapping $\vec{h}\cdot\vec{\sigma}\to\Psi\to\vec{s}$ of Sect. 2, where $\vec{h}\cdot\vec{\sigma}$ is the Hamiltonian, $\Psi$ is the spinor and $\vec{s}$ is the three-dimensional Bloch vector, provides a triple of relevant quantities to characterize an SU(2)SU(2)–based quantum system. The Bloch vector $\vec{s}$ is an observable that visualizes the winding number of the quantum state. To study the Josephson effect we have introduced the periodic order parameter $\Delta=|\Delta|e^{ix/L}$ with the parameter $L$, which induces the spatial winding number $w=l/2\pi L$ of the Bloch vector. $L$ depends on the boundary conditions and can be controlled in the case of the BdG Hamiltonian through the supercurrent as well as through the vector potential of an external electromagnetic field.
An important question regarding the observation of the winding number by using the Bloch vector is whether a closed Bloch-vector trajectory is always accompanied by closed spinor trajectory? The answer is no because this depends on the definition of the spinor. Our loose use of the spinor definition in Eq. (2) or in Eq. (11) indicates that for the latter definition the winding numbers do not agree: The spinor of Eq. (11) has the period $4n\pi L$ but the Bloch vector in Eq. (12) is periodic in $x$ with period $2n\pi L$.
Winding numbers also appear without a periodic order parameter, for instance, in chiral systems. This is briefly mentioned in terms of the 2D Dirac Hamiltonian in Sect. 3.1.1. The main difference is that the winding trajectory of the Bloch vector is parametrized in Fourier space rather than in real space. Moreover, the winding number in this case is $w=1$, which cannot be changed by an external field. But a special unitary transformation can be applied to the Dirac Hamiltonian that creates a phase factor for the Dirac mass which affects the winding number. However, the physical interpretation of the transformed Hamiltonian is not obvious, except for special cases such as the BdG Hamiltonian of a superconductor or a moving polariton condensate [15]. We can summarize that there is a unitary transformation which changes the winding number. This property could be useful once we fully understand how to perform the unitary transformation in a real system.
Finally, it should be emphasized that the Hamiltonians discussed in this work are Hermitian. An extension by including non-Hermitian terms, as used in the concept of “non-Bloch BdG Hamiltonians” [13, 14], might be interesting but should be left for future projects. In particular, to compare the role of edge modes in the Hermitian case with their role in the non-Hermitian case could offer new insights into the effect of the environment on the quantum system. Another extension of the present work could be based on other order parameters. For instance, we can consider piecewise linear phases on short intervals with alternating slopes $\varphi=\pm x/L$, where the phase is still continuous. The matching of the eigenfunctions at the sign-switching points creates additional exponential solutions, creating a special kind of localized wave function. Moreover, if the steps between sign switches are random, this mimics a random phase similar to those in disordered systems.